# Unsupervised Meta-Learning via In-Context Learning

### v1

@@ 6 Conclusion
In this paper, we introduce CAMeLU, a novel approach to UML that leverages the in-context learning capabilities of transformer architectures to extract context from the support samples and make effective predictions on the query data. CAMeLU reframes meta-learning as a sequence modeling problem, where support images provide task context for predicting query images. At the core of CAMeLU is a novel task creation mechanism that generates diverse tasks from an unlabeled dataset, promoting effective generalization to unseen tasks. Our experimental results showcase the superiority of CAMeLU over existing UML methods, highlighting the applicability of the proposed method to domains different from the training one. Notably, CAMeLU can generalize to new domains with a single forward pass (real-time predictions), and it even outperforms its supervised counterpart thanks to its task creation mechanism. Furthermore, the proposed model can be stored and trained with a single GPU with only 4GB of VRAM, underscoring its efficiency in learning-to-learn in-context, rather than using a meta-training phase typical of previous meta-learning approaches.

### v2

@@ 5 Conclusion
In this paper, we introduce CAMeLU, a novel approach for UML that leverages the in-context learning capabilities of transformer architectures to extract context from the support samples and make effective predictions on the query data. CAMeLU reframes meta-learning as a sequence modeling problem, where support images provide task context for predicting query images. At the core of CAMeLU is a novel task creation mechanism that generates diverse tasks from an unlabeled dataset, promoting effective generalization to unseen tasks. Our experimental results showcase the superiority of CAMeLU over existing UML methods, highlighting the applicability of the proposed method to domains different from the training one. Notably, CAMeLU can generalize to new domains with a single forward pass (real-time predictions), and it even outperforms its supervised counterpart thanks to its task creation mechanism. Furthermore, the proposed model can be stored and trained with a single GPU with only 8GB of VRAM, underscoring its efficiency in learning-to-learn in-context, rather than using a meta-training phase typical of previous meta-learning approaches.